OpenStudy (anonymous):

how to find the mid point of the line segment joining (2,3) and (4,5)??

OpenStudy (anonymous):

x= (x sub 1+ x sub 2) all over 2 y= (y sub 1 + y sub 2) all over 2 x= 2+4 all over 2 x= 3 y= 3 + 5 allover 2 y = 4 midpoint (3,4)

OpenStudy (anonymous):

@bii17 is this kinda mean??

OpenStudy (anonymous):



OpenStudy (anonymous):

Thank you guys :)